# DGNS: Deformable Gaussian Splatting and Dynamic Neural Surface for Monocular Dynamic 3D Reconstruction

### v2

@@ 5.3 Ablation study
Surface-aware density control: To illustrate the effect of surface-aware density control, we present two samples of 3D reconstructions with and without surface-aware density control in Fig. 6. The results show that, in the absence of surface-aware density control, a greater number of points are dispersed away from the surface, resulting in a floating appearance. In contrast, applying surface-aware density control, points are more concentrated and closely aligned with the surface, demonstrating improved reconstruction fidelity and surface adherence.
@@
Figure 6: Demonstration of surface-aware density control. The middle image in each sample shows the mesh from the DGS module with surface-aware density control, while the rightmost image shows the result without density control.
Depth Normals CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$
1.006 0.134 34.89
0.831 0.126 35.11
0.697 0.121 35.03
0.502 0.116 35.14
Table 3: Ablation study of depth and normal supervision.
@@
Depth and normal supervision: The ablation study results in Tab 3 highlight the complementary effects of depth and normal information on the quality of 3D reconstruction and novel-view synthesis. When neither depth nor normal data is used, the model exhibits its lowest performance, underscoring the limitations of relying solely on other cues. Adding either depth or normal information individually leads to noticeable improvements, with depth contributing more to spatial alignment accuracy, while normals enhance surface detail and orientation. This suggests that each type of information addresses different aspects of the reconstruction task. When both depth and normal data are combined, the model achieves optimal performance across all metrics, indicating the complementary roles of these inputs. Depth data enhances spatial positioning, while normals provide detailed surface cues, resulting in a more accurate and coherent 3D reconstruction. This improved 3D structure also benefits novel-view synthesis by addressing the challenges of monocular dynamic reconstruction, a highly under-constrained optimization problem. Overall, the results emphasize that while depth and normal information each offer unique benefits, integrating both is crucial for achieving high fidelity in 3D reconstructions and producing realistic novel views.
6 Conclusion
@@
This paper presented our method DGNS, a hybrid framework combining Deformable Gaussian Splatting and Dynamic Neural Surfaces to address the challenges of novel-view synthesis and 3D reconstruction in dynamic scenes. Through surface-aware density control, efficient ray-sampling, and depth supervision, our approach leverages interactions between DGS and DNS to achieve state-of-the-art rendering and geometric accuracy, closely aligning with ground truth and outperforming existing baselines. Experiments on D-NeRF and Dg-mesh datasets demonstrate the robustness of DGNS across complex dynamic scenes, offering a scalable solution for both geometry and appearance modeling.
Limitation: This work primarily addresses the concerns around accuracy as in Fig 1 1, although DGNS incorporates efficient ray-sampling, the DNS module still requires substantial computational resources and longer convergence times compared to DGS, making it the primary speed bottleneck in the framework. Additionally, including both modules during the training stage increases the overall memory footprint, which may limit scalability and efficiency, especially on large datasets or scenes with high complexity.
Future work: Future work will focus on enhancing the computational efficiency of MLP-based DNS modules and extending DGNS’s application to handle more complex dynamic scenes.

### v3

@@ 5.3. Ablation Study
Surface-aware Density Control. To illustrate the effect of surface-aware density control, we present two samples of 3D reconstructions with and without surface-aware density control in Fig. 8. The results show that, in the absence of surface-aware density control, a greater number of points are dispersed away from the surface, resulting in a floating appearance. In contrast, applying surface-aware density control, points are more concentrated and closely aligned with the surface, demonstrating improved reconstruction fidelity and surface adherence.
@@
Figure 8. Demonstration of surface-aware density control. The middle image in each sample shows the mesh from the DGS module with surface-aware density control, while the rightmost image shows the result without density control.
@@
Depth and Normal Supervision. The ablation study results in Tab. 3 highlight the complementary effects of depth and normal information on the quality of 3D reconstruction and novel-view synthesis. The depth map includes well-used $\alpha$-blending and filtered depth map as shown in Eq. (5). When neither depth nor normal data are used, the model exhibits its lowest performance, underscoring the limitations of relying solely on other cues. Adding depth or normal information individually leads to noticeable improvements, with depth contributing more to spatial alignment accuracy, while normals enhance surface detail and orientation. This suggests that each type of information addresses different aspects of the reconstruction task. As for the effect of the depth map, the filtered depth map showed more promising results than $\alpha$-blending depth maps, highlighting the importance of depth filtering. When both filtered depth and normal data are combined, the model achieves optimal performance across all metrics, indicating the complementary roles of these inputs. Depth data enhances spatial positioning, while normals provide detailed surface cues, resulting in a more accurate and coherent 3D reconstruction. This improved 3D structure also benefits novel-view synthesis by addressing the challenges of monocular dynamic reconstruction, a highly under-constrained optimization problem. Overall, the results emphasize that while depth and normal information offer unique benefits, integrating both is crucial for achieving high fidelity in 3D reconstructions and producing realistic novel views.
Table 3. Ablation study of depth and normal supervision.
$\alpha$-Depth Filtered Depth Normals CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ 1.006 0.134 34.89 0.831 0.126 35.11 0.746 0.123 34.98 0.697 0.121 35.02 0.648 0.119 35.06 0.502 0.116 35.14
6. Conclusion
@@
This paper presented our method, DGNS, a hybrid framework combining Deformable Gaussian Splatting and Dynamic Neural Surfaces to address the challenges of novel-view synthesis and 3D reconstruction in dynamic scenes. Through surface-aware density control, efficient ray-sampling, and depth supervision, our approach leverages interactions between DGS and DNS to achieve state-of-the-art rendering and geometric accuracy.Experiments on D-NeRF and Dg-mesh datasets demonstrate the robustness of DGNS across complex dynamic scenes, offering a scalable solution for both geometry and appearance modeling. Limitations: This work primarily addresses the concerns around accuracy as in Fig. 1, although DGNS incorporates efficient ray-sampling, the DNS module still requires substantial computational resources and longer convergence times compared to DGS, making it the primary speed bottleneck in the framework. Additionally, including both modules during the training stage increases the overall memory footprint, which may limit scalability and efficiency, especially on large datasets or scenes with high complexity. Future work will focus on enhancing the computational efficiency of MLP-based DNS modules and extending DGNS’s application to handle more complex dynamic scenes.